OpenStudy (anonymous):

James left a bin outside in his garden to collect rainwater. He notices that 1 over 4 gallon of water fills 2 over 3 of the bin. Write and solve an equation to find the amount of water that will fill the entire bin.

OpenStudy (anonymous):

1/4 / 2/3 = x            1/4 * 3/2 = x  3/8 = x My answer . Am I on the right track?

OpenStudy (anonymous):

@jim_thompson5910 am I the right track here

jimthompson5910 (jim_thompson5910):

Yeah you'll start off with